# Explainable Multi-modal Time Series Prediction with LLM-in-the-Loop

### v2

@@ 4.1 Experimental Setup
Baselines, Evaluation Metrics, and Setup We compare TimeXL with state-of-the-art baselines for time series prediction, including Autoformer [94], Dlinear [93], Crossformer [95], TimesNet [96], PatchTST [5], iTransformer [97], FreTS [100], TSMixer [98], TimeMixer [99] (classification implemented by TSlib [96]) and LLM-based methods like LLMTime [101], PromptCast [79], OFA [66], Time-LLM [69], TimeCMA [72] and FSCA [73], where LLMTime and PromptCast don’t need fine-tuning. While some methods are primarily used for time series prediction with continuous values, they can be easily adapted for discrete value prediction. We also evaluate the multi-modal time series methods. Besides the Time-LLM, TimeCMA and FSCA where input text is used for embedding reprogramming and alignment, we also evaluate Multi-modal PatchTST and Multi-modal iTransformer from [47], as well as TimeCAP [81]. We evaluate the discrete prediction via F1 score and AUROC (AUC) score, due to label imbalance in real-world time series datasets. All datasets are split 6:2:2 for training, validation, and testing. All results are averaged over five experimental runs. We alternate different embedding methods for texts based on its average length (Sentence-BERT [92] for finance dataset and BERT [91] for the others). More details are provided in Appendix A.2A.3A.4.
@@ 4.2 Performance Evaluation
The results of predictive performance are shown in Table 1. It is notable that multi-modal methods generally outperform time series methods across all datasets. These methods include LLM methods (e.g., Time-LLM [69], TimeCMA [72], FSCA [73]) that leverage text embeddings to enhance time series predictions. Moreover, the multi-modal variants (MM-iTransformer and MM-PatchTST [47]) improve the performance of state-of-the-art time series methods, suggesting the benefits of integrating real-world contextual information. Besides, TimeCAP [81] integrates the predictions from both modalities, further improving the predictive performance. TimeXL constantly achieves the highest F1 and AUC scores, consistently surpassing both time series and multi-modal baselines by up to 8.9% of AUC (compared to TimeCAP on Weather dataset). This underscores the advantage of TimeXL, which synergizes multi-modal time series encoder with language agents to enhance interpretability and thus predictive performance in multi-modal time series.
@@ 4.3 Explainable Multi-modal Prototypes
Next, we present the explainable multi-modal prototypes rendered by TimeXL, which establishes the case-based reasoning process. Figure 3 shows a subset of time series and text prototypes learned on the weather dataset. The time series prototypes demonstrate the typical temporal patterns aligned with different real-world weather conditions (i.e., rain and not rain). For example, a constant or decreasing humidity at a moderate level, combined with high and steady air pressure, typically indicates a non-rainy scenario. The consistent wind direction is also a sign of mild weather conditions. On the contrary, high humidity, low and fluctuating pressure, along with variable winds typically reveal an unstable weather system ahead. In addition to time series, the text prototypes also highlight consistent semantic patterns for different weather conditions, such as the channel-specific (e.g., drier air moving into the area, strengthening of high-pressure system) and overall (e.g., a likelihood of dry weather) descriptions of weather activities. In Appendix C.1, we also present more multi-modal prototypes for the weather dataset in Figure 9, for the finance dataset in Figure 10, and for healthcare datasets in Figures 11 and 12. TimeXL provides coherent and informative prototypes from the exploitation of time series and its real-world contexts, which facilitates both prediction and explanation.
Figure 3: Key time series prototypes and text prototypes learned on weather data. Each row in the figure represents a time series prototype with different channels.
Figure 4: Multi-modal case-based reasoning example on weather data. The left part illustrates the reasoning process for both the original and refined text in TimeXL, with matched prototype-input pairs highlighted in the same color along with their similarity scores. The right part presents the time series reasoning in TimeXL, where matched prototypes are overlaid on the time series.
@@ 4.4 Multi-modal Case-based Reasoning
Building upon the multi-modal prototypes, we present a case study on the testing set of weather data, comparing the original and TimeXL’s reasoning processes to highlight its explanatory capability, as shown in Figure 4. In this case, the original text is incorrectly predicted as not rain. We have three key observations: (1) The refinement process filters the original text to emphasize weather conditions more indicative of rain, guided by reflections from training examples. The refined text preserves the statement on stability while placing more emphasis on humidity and wind as key indicators. (2) Accordingly, the matched segment-prototype pairs from the original text focus more on temperature stability and typical diurnal variations, while the matched pairs in the refined text highlights wind variability, moisture transport, and approaching weather system, aligning more with rain conditions. (3) Furthermore, the reasoning on time series provides a complementary view for assessing weather conditions. The matched time series prototypes identify high humidity and its drop-and-rise trends, wind speed fluctuations and directional shifts, and the declining phase of air pressure fluctuations, all of which are linked to the upcoming rainy conditions. The matched multi-modal prototypes from TimeXL demonstrate its effectiveness in capturing relevant information for both predictive and explanatory analysis. We also provide a case study on finance data in Figure 13, where textual explanations are generated at the granularity of a half-sentence.
@@ 4.5 Iterative Analysis
Figure 5: Iterative analysis: the text quality and TimeXL performance over iterations.
@@ 4.6 Ablation Studies
Table 2: Ablation studies on real-world multi-modal time series datasets, evaluated by F1 score.
Ablation Variants Weather Finance TP MT
Encoder Time Series 0.602 0.585 0.889 0.861
Text 0.567 0.472 0.871 0.840
Multi-modal 0.674 0.619 0.934 0.937
LLM Text 0.645 0.496 0.974 0.901
Text + Prototype 0.667 0.544 0.987 0.952
Fusion Select-Best 0.674 0.619 0.987 0.952
TimeXL 0.696 0.631 0.987 0.956
In this subsection, we present the component ablations of TimeXL, as shown in Table 2, where the texts are refined using the reflections with the best validation performance. Firstly, the multi-modal encoder consistently outperforms single-modality variants across all datasets, highlighting the benefit of integrating time series and textual modalities. Secondly, the prediction LLM exhibits better performance than text-only encoder, underscoring the advantage of LLM’s contextual understanding and reasoning. Furthermore, the text prototypes consistently improve the predictive performance of LLM, underscoring the effectiveness of explainable artifacts from the multi-modal encoder, in terms of providing relevant contextual guidance. Finally, the fusion of prediction LLM and multi-modal encoder further boosts the predictive performance that surpasses the best of both multi-modal encoder and prediction LLM. These observations demonstrate the advantage of our framework synergizing the time series model and LLM for mutually augmented prediction. In Appendix B, full results (F1 and AUC) of TimeXL component ablation are provided in Table 4, other ablations and studies (base LLM, quantitative evaluation of case-based explanations, learning objectives) are provided in Table 5 and Figures 678.
@@ A.1 Dataset Statistics
In this subsection, we provide more details of the real-world datasets we used for the experiments. The data statistics are summarized in Table 3, including the meta information (e.g., domain resolution, duration of real-world time series records), the number of channels and timesteps and so on. We used the weather and healthcare datasets in TimeCAP [81], and the finance dataset in [13].
@@
The Finance dataset contains the daily record of the raw material prices together with 14 related indices ranging from January 2017 to July 2024222https://www.indexmundi.com/commodities, with news articles gathered from S&P Global Commodity Insights. The task is to predict if future prices will increase by more than 1%, decrease by more than 1%, or exhibit a neutral trend on the next business day, given the last 5 business days of stock price data and news.
@@ A.4 Environment
Appendix B More Ablation Studies
F1 AUC F1 AUC F1 AUC F1 AUC
Encoder Time Series 0.602 0.691 0.585 0.751 0.889 0.957 0.861 0.965
Text 0.567 0.658 0.472 0.636 0.871 0.941 0.840 0.887
Multi-modal 0.674 0.767 0.619 0.791 0.934 0.974 0.937 0.988
LLM Text 0.645 0.724 0.496 0.627 0.974 0.967 0.901 0.969
Text + Prototype 0.667 0.739 0.544 0.662 0.987 0.983 0.952 0.976
Fusion Select-Best 0.674 0.767 0.619 0.791 0.987 0.983 0.952 0.988
TimeXL 0.696 0.808 0.631 0.797 0.987 0.996 0.956 0.997
Component ablations of TimeXL: Here we present the full results of TimeXL component ablations in Table 4. In addition to F1 scores, the results of AUC scores consistently demonstrate the importance of contextual information, the effectiveness of prototype from the multi-modal encoder, as well as the advantage of prediction fusion.
@@ Appendix B More Ablation Studies
Effect of case-based explanations on LLM predictions and other quality assessments: Moreover, we assess how the number of allocated case-based explanations enhances the prediction LLM, as shown in Figure 7. We conduct experiments on weather and finance datasets, demonstrating that incorporating more relevant case-based explanations consistently improves prediction performance. This trend highlights the effectiveness of explainable artifacts in providing meaningful contextual guidance. Importantly, this analysis also serves as a proxy for quantitatively evaluating explanation quality, where retrieving higher-quality case-based examples tends to yield greater performance gains. Beyond this quantitative assessment, alternative strategies can be employed to evaluate explanation quality more explicitly, for example, using LLM as a judge [103] to rate the relevance and helpfulness of explanations in a human-understandable manner, with respect to ground truth.
@@
Effect of different base LLMs: To explore the effect of different base LLMs (Gemini-2.0 Flash and GPT-4o-mini), we provide the experiment results and analysis on Healthcare (Test-Positive) data. For fair comparisons, we used the same prompts and ran the same number of iterations and followed the same settings detailed in Appendix A.2,  A.3, and  A.4. The results are shown in Table 5. It can be observed that GPT-4o clearly outperforms both GPT-4o-mini and Gemini-2.0-Flash, due to its better reasoning and contextual understanding capabilities (larger model size & pre-training corpus). Both GPT-4o-mini and Gemini-2.0-Flash are still competitive compared with baselines listed in Table 1. It reveals the impact of the LLM capability on the effectiveness of our framework, especially in tasks demanding real-world context understanding. We also provide the plot of iterative analysis in Figure 8, where we can also observe the performance improvement over iterations for different base LLMs.
Figure 6: Ablation study of TimeXL encoder learning objectives.
@@
Figure 8: Iterative study of different base LLMs on Healthcare (Test-Positive) dataset.
Appendix C Explainable Multi-modal Prototypes and Case Study
C.1 Multi-modal Prototypes for All Datasets
@@
C.2 Case-based Reasoning Example on Finance Dataset
We provide another case-based reasoning example to demonstrate the effectiveness of TimeXL in explanatory analysis, as shown in Figure 13. In this example, the original text is incorrectly predicted as neutral instead of a decreasing trend of raw material prices. We have a few key observations based on the results. First, the refinement LLM filters the original text to emphasize economic and market conditions more indicative of a declining trend, based on the reflections from training examples. The refined text preserves discussions on port inventories and steel margins while placing more emphasis on subdued demand, thin profit margins, and bearish market sentiment as key indicator of prediction. Accordingly, the case-based explanations from the original text focus more on inventory management and short-term stable patterns, while those in the refined text highlight demand contraction, production constraints, and macroeconomic uncertainty, which is more consistent with a decreasing trend. Furthermore, the reasoning on time series provides a complementary view for predicting iron ore price trends. The time series explanations identify declining price movements across multiple indices. In general, the multi-modal explanations based on matched prototypes from TimeXL demonstrate its effectiveness in capturing relevant iron ore market condition for both predictive and explanatory analysis.
Figure 9: More multi-modal prototypes learned from Weather dataset. Each row in the figure represents a time series prototype.
Figure 10: Key multi-modal prototypes learned from Finance dataset. Each row in the figure represents a time series prototype.
Figure 11: Key multi-modal prototypes learned from Healthcare (Test-positive) dataset. Each row in the figure represents a time series prototype.
Figure 12: Key multi-modal prototypes learned from Healthcare (Mortality) dataset. Each row in the figure represents a time series prototype.
Figure 13: Multi-modal case-based reasoning example on Finance dataset.
Appendix D Designed Prompts for Experiments

### v3

@@ 4.1 Experimental Setup
Baselines, Evaluation Metrics, and Setup We compare TimeXL with state-of-the-art baselines for time series prediction, including Autoformer wu2021autoformer , Dlinear zeng2023transformers , Crossformer zhang2023crossformer , TimesNet wutimesnet , PatchTST nietime , iTransformer liuitransformer , FreTS yi2024frequency , TSMixer chentsmixer , TimeMixer wangtimemixer (classification implemented by TSlib wutimesnet ) and LLM-based methods like LLMTime gruver2023llmtime , PromptCast xue2023promptcast , OFA zhou2023one , Time-LLM jintime , TimeCMA liu2024timecma and FSCA hu2025contextalignment , where LLMTime and PromptCast don’t need fine-tuning. While some methods are primarily used for time series prediction with continuous values, they can be easily adapted for discrete value prediction. We also evaluate the multi-modal time series methods. Besides the Time-LLM, TimeCMA and FSCA where input text is used for embedding reprogramming and alignment, we also evaluate Multi-modal PatchTST and Multi-modal iTransformer from liu2024time , as well as TimeCAP lee2024timecap . We evaluate the discrete prediction via F1 score and AUROC (AUC) score, due to label imbalance in real-world time series datasets. All datasets are split 6:2:2 for training, validation, and testing. All results are averaged over five experimental runs. We alternate different embedding methods for texts based on its average length (Sentence-BERT reimers-2019-sentence-bert for finance dataset and BERT kenton2019bert for the others). More details are provided in Appendix A.2A.3A.4.
@@ 4.2 Performance Evaluation
The results of predictive performance are shown in Table 1. It is notable that multi-modal methods generally outperform time series methods across all datasets. These methods include LLM methods (e.g., Time-LLM jintime , TimeCMA liu2024timecma , FSCA hu2025contextalignment ) that leverage text embeddings to enhance time series predictions. Moreover, the multi-modal variants (MM-iTransformer and MM-PatchTST liu2024time ) improve the performance of state-of-the-art time series methods, suggesting the benefits of integrating real-world contextual information. Besides, TimeCAP lee2024timecap integrates the predictions from both modalities, further improving the predictive performance. TimeXL constantly achieves the highest F1 and AUC scores, consistently surpassing both time series and multi-modal baselines by up to 8.9% of AUC (compared to TimeCAP on the Weather dataset). This underscores the advantage of TimeXL, which synergizes multi-modal encoder with language agents to enhance interpretability and thus predictive performance in multi-modal time series.
Figure 3: Key time series prototypes and text prototypes learned on the Weather dataset. Each row in the figure represents a time series prototype with different channels.
@@ 4.3 Explainable Multi-modal Prototypes
Next, we present the explainable multi-modal prototypes rendered by TimeXL, which establishes the case-based reasoning process. Figure 3 shows a subset of time series and text prototypes learned on the weather dataset. The time series prototypes demonstrate the typical temporal patterns aligned with different real-world weather conditions (i.e., rain and not rain). For example, a constant or decreasing humidity at a moderate level, combined with high and steady air pressure, typically indicates a non-rainy scenario. The consistent wind direction is also a sign of mild weather conditions. On the contrary, high humidity, low and fluctuating pressure, along with variable winds typically reveal an unstable weather system ahead. In addition to time series, the text prototypes also highlight consistent semantic patterns for different weather conditions, such as the channel-specific (e.g., drier air moving into the area, strengthening of high-pressure system) and overall (e.g., a likelihood of dry weather) descriptions of weather activities. In Appendix D.1, we also present more multi-modal prototypes for the weather dataset in Figure 9, for the finance dataset in Figure 10, and for healthcare datasets in Figures 11 and 12. TimeXL provides coherent and informative prototypes from the exploitation of time series and its real-world contexts, which facilitates both prediction and explanation.
Figure 4: Multi-modal case-based reasoning example on the Weather dataset. The left part illustrates the reasoning process for both the original and refined text in TimeXL, with matched prototype-input pairs highlighted in the same color along with their similarity scores. The right part presents the time series reasoning in TimeXL, where matched prototypes are overlaid on the time series.
@@ 4.4 Multi-modal Case-based Reasoning
Building upon the multi-modal prototypes, we present a case study on the testing set of weather data, comparing the original and TimeXL’s reasoning processes to highlight its explanatory capability, as shown in Figure 4. In this case, the original text is incorrectly predicted as not rain. We have three key observations: (1) The refinement process filters the original text to emphasize weather conditions more indicative of rain, guided by reflections from training examples. The refined text preserves the statement on stability while placing more emphasis on humidity and wind as key indicators. (2) Accordingly, the matched segment-prototype pairs from the original text focus more on temperature stability and typical diurnal variations, while the matched pairs in the refined text highlights wind variability, moisture transport, and approaching weather system, aligning more with rain conditions. (3) Furthermore, the reasoning on time series provides a complementary view for assessing weather conditions. The matched time series prototypes identify high humidity and its drop-and-rise trends, wind speed fluctuations and directional shifts, and the declining phase of air pressure fluctuations, all of which are linked to the upcoming rainy conditions. The matched multi-modal prototypes from TimeXL demonstrate its effectiveness in capturing relevant information for both predictive and explanatory analysis. We also provide a case study on the Finance dataset in Figure 13, where textual explanations are generated at the granularity of a half-sentence.
@@ 4.5 Iterative Analysis
Figure 5: The text quality and TimeXL performance over iterations.
@@ 4.6 Ablation Studies
F1 AUC F1 AUC F1 AUC F1 AUC
Encoder Time Series 0.602 0.691 0.585 0.751 0.889 0.957 0.861 0.965
Text 0.567 0.658 0.472 0.636 0.871 0.941 0.840 0.887
Multi-modal 0.674 0.767 0.619 0.791 0.934 0.974 0.937 0.988
LLM Text 0.645 0.724 0.496 0.627 0.974 0.967 0.901 0.969
Text + Prototype 0.667 0.739 0.544 0.662 0.987 0.983 0.952 0.976
Fusion Select-Best 0.674 0.767 0.619 0.791 0.987 0.983 0.952 0.988
TimeXL 0.696 0.808 0.631 0.797 0.987 0.996 0.956 0.997
We present the component ablations of TimeXL in Table 2, where the texts are refined using the reflections with the best validation performance. Firstly, the multi-modal encoder consistently outperforms single-modality variants across all datasets, highlighting the benefit of integrating time series and textual modalities. Secondly, the prediction LLM exhibits better performance than text-only encoder, underscoring the advantage of LLM’s contextual understanding and reasoning. Furthermore, the text prototypes consistently improve the predictive performance of LLM, underscoring the effectiveness of explainable artifacts from the multi-modal encoder, in terms of providing relevant contextual guidance. Finally, the fusion of prediction LLM and multi-modal encoder further enhances the predictive performance, surpassing the best results achieved by either component alone. These observations demonstrate the advantage of our framework synergizing the time series model and LLM for mutually augmented prediction. In Appendix B, additional ablation and model studies are presented, including analyses of the multi-modal encoder (Figure 6; Table 466), quantitative evaluation of case-based explanations (Figure 7), and comparisons of base LLMs (Table 7; Figure 8).
@@ A.1 Dataset Statistics
In this subsection, we provide more details of the real-world datasets we used for the experiments. The data statistics are summarized in Table 3, including the meta information (e.g., domain resolution, duration of real-world time series records), the number of channels and timesteps and so on. We use the Weather and Healthcare datasets in TimeCAP [65], and a Finance dataset extended from [12].
@@
The Finance dataset contains the daily record of the raw material prices together with 14 related indices ranging from January 2017 to July 2024222https://www.indexmundi.com/commodities with news articles. The task is to predict if future prices will increase by more than 1%, decrease by more than 1%, or exhibit a neutral trend on the next business day, given the last 5 business days of stock price data and news.
@@ A.4 Environment
Appendix B More Ablation Studies and Model Analysis
@@
Figure 6: Ablation study of multi-modal encoder learning objectives.
Sensitivity of regularization strengths: To quantitatively evaluate the sensitivity of regularization strengths in the learning objectives, we report the performance of the multi-modal encoder on the Weather dataset (using the refined test texts), as detailed in Appendix A.2. As shown in Table 4, model performance remains relatively stable across different values. For both modalities, moderate values of $\lambda_{1}$ and $\lambda_{2}$ tend to yield better performance, suggesting the clustering and evidencing effectively guide meaningful prototype learning. A slight diversity constraint ($\lambda_{3}$) also helps to keep a compact interpretation prototype space.
Table 4: Sensitivity analysis of different $\lambda$ values for multi-modal encoder on the Weather dataset.
Modality 0.1 0.15 0.2 0.25 0.3
[…]
Sensitivity of prototype lengths: We evaluate the sensitivity of the multi-modal encoder to prototype length by varying the single-kernel sizes for both time-series and text modalities on the Weather and Finance datasets (using the refined test texts), as shown in Table 6 and 6. We observe that performance is relatively stable across a range of prototype lengths on the Weather dataset, with optimal results typically achieved at moderate lengths for both modalities. For time series, the 8-hour segments are often sufficient to capture the typical temporal patterns of weather conditions, and a 12-token text window (with BERT) provides enough local contexts for weather prediction. For the Finance dataset, which operates on a business-day granularity, performance improves with longer lengths of time series prototypes that capture important financial trends. For the textual modality, we use Sentence-BERT to embed financial news at the half-sentence granularity, and shorter text windows (1-2 segments) typically provide compact and indicative market conditions. These results suggest that our model maintains robust to prototype length selection across a reasonable range, guided by the temporal characteristics of each dataset.
Table 5: Sensitivity analysis of prototype lengths for multi-modal encoder on the Weather dataset.
Modality Length F1 AUC
[…]
Table 6: Sensitivity analysis of prototype lengths for multi-modal encoder on the Finance dataset.
Modality Length F1 AUC
[…]
Effect of case-based explanations on LLM predictions and other quality assessments: Moreover, we assess how the number of allocated case-based explanations enhances the prediction LLM, as shown in Figure 7. We conduct experiments on the Weather and Finance datasets, demonstrating that incorporating more relevant case-based explanations consistently improves prediction performance. This trend highlights the effectiveness of explainable artifacts in providing meaningful contextual guidance. Importantly, this analysis also serves as a proxy for quantitatively evaluating explanation quality, where retrieving higher-quality case-based examples tends to yield greater performance gains. Beyond this quantitative assessment, alternative strategies can be employed to evaluate explanation quality more explicitly, for example, using LLM as a judge [85] to rate the relevance and helpfulness of explanations in a human-understandable manner, with respect to ground truth.
@@
Effect of different base LLMs: To explore the effect of different base LLMs (Gemini-2.0 Flash and GPT-4o-mini), we provide the experiment results and analysis on the Healthcare (Test-Positive) dataset. For fair comparisons, we used the same prompts and ran the same number of iterations and followed the same settings detailed in Appendix A.2,  A.3, and  A.4. The results are shown in Table 7. It can be observed that GPT-4o clearly outperforms both GPT-4o-mini and Gemini-2.0-Flash, due to its better reasoning and contextual understanding capabilities (larger model size & pre-training corpus). Both GPT-4o-mini and Gemini-2.0-Flash are still competitive compared with baselines listed in Table 1. It reveals the impact of the LLM capability on the effectiveness of our framework, especially in tasks demanding real-world context understanding. We also provide the plot of iterative analysis in Figure 8, where we can also observe the performance improvement over iterations for different base LLMs.
@@
Figure 8: Iterative study of different base LLMs on the Healthcare (Test-Positive) dataset.
Appendix C Computation Cost for Iterative Optimization and Inference
We quantitatively analyze the token usage and runtime of the iterative process. Following the prompts and strategy detailed in Appendix E, each iteration over 4000 training and validation samples (average length of 400 tokens, around 300 words) requires about 9.74 M input tokens and yield 1.64 M output tokens. With GPT‑4o (with input latency as TTFT $\approx$ 0.45 seconds and output speed $\approx$ 156 tokens/second), it corresponds to around 3.9 hours serially per iteration (prediction takes approximately 0.5 hours, reflection 0.08 hours, and refinement 3.35 hours). As prediction and refinement steps can also be parallelized, we can reduce the iteration time to about 12 minutes with 20 concurrent calls. Using a more efficient LLM base model Gemini-Flash-2.0 with TTFT $\approx$ 0.3 seconds and $\approx$ 236 output tokens/second, the iteration time can be further reduced to 8 minutes.
In the testing stage, we apply the fixed reflection selected during validation to refine the test texts. Based on our prompt lengths, the refinement and prediction steps take approximately 3.46 seconds per sample using GPT‑4o, and 2.29 seconds per sample for Gemini-2.0-Flash, which is reasonable for deployment in real-world applications.
In addition to efficient LLM backbones and parallelized querying, we discuss further strategies to lower computational cost. As shown in Figure 5, 1-2 iterations often yield clear performance gains. In practice, the iteration loop can be adaptively controlled by monitoring validation improvements and stopping early once gains fall below a predefined threshold. To further reduce the cost, we can also do selective refinement by skipping training examples that were already predicted accurately in previous iterations. These strategies help make our method more computationally tractable.
Appendix D Explainable Multi-modal Prototypes and Case Study
D.1 Multi-modal Prototypes for All Datasets
@@
D.2 Case-based Reasoning Example on the Finance Dataset
We provide another case-based reasoning example to demonstrate the effectiveness of TimeXL in explanatory analysis, as shown in Figure 13. In this example, the original text is incorrectly predicted as neutral instead of a decreasing trend of raw material prices. We have a few key observations based on the results. First, the refinement LLM filters the original text to emphasize economic and market conditions more indicative of a declining trend, based on the reflections from training examples. The refined text preserves discussions on port inventories and steel margins while placing more emphasis on subdued demand, thin profit margins, and bearish market sentiment as key indicator of prediction. Accordingly, the case-based explanations from the original text focus more on inventory management and short-term stable patterns, while those in the refined text highlight demand contraction, production constraints, and macroeconomic uncertainty, which is more consistent with a decreasing trend. Furthermore, the reasoning on time series provides a complementary view for predicting raw material price trends. The time series explanations identify declining price movements across multiple indices. In general, the multi-modal explanations based on matched prototypes from TimeXL demonstrate its effectiveness in capturing relevant raw material market condition for both predictive and explanatory analysis.
Figure 9: More multi-modal prototypes learned from the Weather dataset. Each row in the figure represents a time series prototype.
Figure 10: Key multi-modal prototypes learned from the Finance dataset. Each row in the figure represents a time series prototype.
Figure 11: Key multi-modal prototypes learned from the Healthcare (Test-Positive) dataset. Each row in the figure represents a time series prototype.
Figure 12: Key multi-modal prototypes learned from the Healthcare (Mortality) dataset. Each row in the figure represents a time series prototype.
Figure 13: Multi-modal case-based reasoning example on the Finance dataset.
Appendix E Designed Prompts for Experiments